EXHIBIT 7.5 Selected Financial Data for Comparable Companies (in millions of USD, except percentages, ratios, and per share dUsing betas in Exhibit 7.5 and the market excess return (market return less risk free rate) of 8%, what is the difference of Supervalu expected return and Whole Foods Market expected return? Enter percentage, round to 2 decimal places.
Answer #1

Beta of Supervalu =1.00

Expected Return of Supervalu=Expected Return of Market , because Beta=1

CAPM Equation

Rs=Rf+Beta*(Rm-Rf)

Rs=Supervalu Return

Rf=Risk Free Rate

Beta =1

Rm=Expected Market Return=Rf+8%

Rs=Rf+1*(Rm-Rf)=Rf+Rm-Rf=Rm=Rf+8%

Rw=Whole Foods Market expected Return

Rw=Rf+Beta*(Rm-Rf)

Beta for WholeFoods Market=0.7

Rw=Rf+0.7*(Rf+8-Rf)

Rw=Rf+0.7*8=Rf+5.6%

Rs=Rf+8%

Difference of Supervalu expected return and Whole Foods Market expected return=Rs-Rw=Rf+8%-(Rf+5.6%)=2.40%

Difference of Supervalu expected return and Whole Foods Market expected return=2.4%